# Using negative controls to identify causal effects with invalid instrumental variables

### v4

@@ Abstract
Many proposals for the identification of causal effects require an instrumental variable that satisfies strong, untestable unconfoundedness and exclusion restriction assumptions. In this paper, we show how one can potentially identify causal effects under violations of these assumptions by harnessing a negative control population or outcome. This strategy allows one to leverage sup-populations for whom the exposure is degenerate, and requires that the instrument-outcome association satisfies a certain parallel trend condition. We develop the semiparametric efficiency theory for a general instrumental variable model, and obtain a multiply robust, locally efficient estimator of the average treatment effect in the treated. The utility of the estimators is demonstrated in simulation studies and an analysis of the Life Span Study.
@@ 1 Introduction
Alternatively, a negative control outcome is a variable for which it is implausible that there can be a causal effect of the exposure. Under the core instrumental variable assumptions, the instrument would then not be associated with the negative control outcome. For example, one might choose a clinical outcome unlikely to be affected by the medication. There is an existing literature on using negative control outcomes as a means to detect and remove bias in an analysis based on confounding adjustment (Lipsitch et al.,, 2010; Sofer et al.,, 2016); however, they have been used less for assessing the instrumental variable assumptions. Davies et al., (2017) describe how negative control outcomes can be used to compare the plausibility of instrumental variable approaches with analyses that assume no unmeasured confounding; Sanderson et al., (2021) use these variables to check for violations of unconfoundedness due to population stratification in Mendelian Randomisation studies.
We go beyond falsification testing and propose conditions for identification and estimation of effects with an invalid instrument variable. Our results rely on a condition on the stability of the association between the instrument and outcome, which is closely related to the parallel trends assumption in the difference-in-difference literature. Intuitively, our assumption states that the instrument-outcome association due to violations of IV.2 and IV.3 in the naïve instrumental variable analysis must equal the instrument-outcome association in a reference population. If our assumption holds, the potential instrument does not necessarily have to satisfy the conditions IV.2 and IV.3, but still must be predictive of the exposure, therefore satisfying IV.1. A special feature of our framework, relative to other negative control strategies, is that it allows one to leverage sub-populations who were not eligible to be exposed. We hope that our methodology may be generally useful in epidemiologic applications where such sub-populations arise.
@@ 2.1 Bias detection using reference populations
Reference population: Let $\mathcal{C}_{0}$ ($\mathcal{C}_{1}$) denote the support of $\mathbf{C}$ in those with $S=0$ ($S=1$). Then $\forall z\in\{0,1\}$ and $\forall\mathbf{c}\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$,
@@
A negative control population where the exposure is not necessarily degenerate but is hypothesised to have no effect, could also form a reference population. However, the concept of a reference population allows for the possibility that the treatment causally impacts the outcome in those with $S=0$, if these individuals were not excluded from access to treatment. In general, it is advantageous to choose a reference population that is similar in important baseline characteristics to the $S=1$ population. This is because to eventually identify a causal effect, we will need to transport associations across from the reference to the $S=1$ population. We emphasise that it is not obvious how other negative control-type designs can generally leverage sub-populations for whom treatment is degenerate. These designs usually involve contrasts between levels of $A$ that are expected to be zero when there is no bias. Hence some variation in exposure is typically required (Lipsitch et al.,, 2010).
$Z$$A$$C$$Y$$U$(a) Target population ($S=1$)
@@
See B.1 of the Appendix for a proof. This statement provides the basis of a valid test of IV.2 and IV.3 in the reference population. To draw conclusions about the target population, we would furthermore need to assume that $\forall z\in\{0,1\}$ and $\forall\mathbf{c}\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$,
@@
In other words, $Z$ is an instrument in $S=0$ if it is an instrument in $S=1$. Then (1) can be used as the basis of a valid falsification test for the IV assumptions in the target population. If we strengthen (2.1) to an ‘iff’ statement and Proposition 2.0.1 to an ‘iff’ result (under a faithfulness-type condition), then we can construct a consistent test. We note that the proposition allows for the distribution of $\mathbf{C}$ to differ between populations, but if the support of $\mathbf{C}$ differs between the standard and reference populations, one may not be able to draw conclusions for all individuals with $S=1$.
@@ 2.2 Point identification
We will focus on identification of the conditional average treatment effect on the treated in the target population:
@@
We note first that Assumption 2 is standard in causal inference. Assumption 3 demands that $Z$ is predictive of the exposure; although this assumption is similar to IV.1, $Z$ is not required to satisfy either IV.2 or IV.3. In particular, our assumptions also allow for $Z$ to be a standard confounder with a direct effect on the outcome (violating the exclusion restriction). When IV.2 and IV.3 hold in both populations, we note that Assumption 4 then automatically also holds. This is because one can show that under the IV conditions, $E(Y^{0}|Z=1,S=s,\mathbf{C}=\mathbf{c})-E(Y^{0}|Z=0,S=s,\mathbf{C}=\mathbf{c})=0$ $\forall\mathbf{c}\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$ and $s=0,1$. Assumption 4 requires that the (conditional) additive association between the potential outcome $Y_{0}$ and $Z$ transports between the negative control and target populations. It is not generally testable. We discuss the plausibility of this assumption below, which is weaker than the full population exchangeability assumption made in the transportability literature (Dahabreh et al.,, 2019).
@@
Our assumptions are stated conditional on covariates $\mathbf{C}$. To ensure Assumption 4 holds, we require that $\mathbf{C}$ includes all effect modifiers of the additive association between $Z$ and $Y^{0}$ that differ in distribution between populations. In terms of design implications, this suggests choosing a reference population that is similar to the target population in terms of characteristics that may modify the association between $Z$ and $Y^{0}$. As argued above, it is also advantageous to incorporate measured confounders for the effect of $Z$ on $Y$, as well as for the effect of $A$ on $Y$, since this may weaken the restrictions on $U$.
@@ 2.3 Negative control outcomes
One can also harness measurements of a negative control outcome in order to generate a reference population. Suppose that there is now only data on individuals from the target population, such that $S=1$ for everyone (we omit this from the conditioning statement to simplify notation). Further, suppose that one collects a random variable $W$ in the target population that satisfies the following assumption
@@ 3.1 Semiparametric Theory
$\Omega=\{\phi(\mathbf{Z},S,\mathbf{C}):E\{\phi(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}=E\{\phi(\mathbf{Z},S,\mathbf{C})|S,\mathbf{C}\}=0\}$
@@ 3.2 De-biased machine learning and multiply robust estimation
The estimation strategies previously described are not generally feasible, because the estimating equations for $\psi^{\dagger}$ involve nuisance parameters that are typically unknown. One approach to deal with this is to plug-in estimates obtained from nonparametric estimators or flexible statistical learning methods (Chernozhukov et al.,, 2018). In A.6 of the Appendix, we therefore use Theorem 3.1 to construct cross-fit de-biased machine learning-based estimators of $\psi^{\dagger}$, and describe sandwich estimators of the asymptotic variance. Nevertheless, to develop a more nuanced understanding of the different bias properties of RAL estimators of $\psi^{\dagger}$, we will instead consider a scenario where parametric working models are used for the conditional expectations/densities that arise in estimating the causal effect of interest. We hence conceptualise bias as potentially arising from parametric model misspecification. The parametric approach will also suggest how nuisance functions should be estimated nonparametrically for optimal performance in terms of their convergence rates.
@@
  • $\mathcal{M}_{3}$: $f(S|\mathbf{Z}=0,\mathbf{C})=f(S|\mathbf{Z}=0,\mathbf{C};\alpha^{\dagger})$, $OR(\mathbf{Z},S,\mathbf{C})=OR(\mathbf{Z},S,\mathbf{C};\rho^{\dagger})$, and $b_{1}(\mathbf{C})=b_{1}(\mathbf{C};\theta^{\dagger}_{1})$.
  • $\mathcal{M}_{4}$: $f(\mathbf{Z}|S=0,\mathbf{C})=f(\mathbf{Z}|S=0,\mathbf{C};\tau^{\dagger})$, $f(S|\mathbf{Z}=0,\mathbf{C})=f(S|\mathbf{Z}=0,\mathbf{C};\alpha^{\dagger})$ and $OR(\mathbf{Z},S,\mathbf{C})=OR(\mathbf{Z},S,\mathbf{C};\rho^{\dagger})$.
At one extreme, $\mathcal{M}_{1}$ requires a correct conditional mean model for outcome after removing the treatment effect: $E\{Y-\beta(\mathbf{A,C})S|\mathbf{Z},S,\mathbf{C}\}$. At the other, $\mathcal{M}_{4}$ requires a correct model for the joint conditional density $f(\mathbf{Z},S|\mathbf{C})$. However, we allow for certain combinations of restrictions on these laws. If our model $f(S|\mathbf{Z},\mathbf{C})$ is correct, then under $\mathcal{M}_{3}$, assumptions on the impact of $\mathbf{Z}$ on the conditional mean of $Y$ given $\mathbf{C}$ in the reference population can be relaxed.
@@ 4 Data analysis
Varying the fitted models in this way enabled us to check the robustness of the estimators $\hat{\psi}_{TSLS}$, $\hat{\psi}_{g-Z}$, $\hat{\psi}_{g-S}$, $\hat{\psi}_{IPW}$, $\hat{\psi}_{MR}$ and $\hat{\psi}_{MR-eff}$ to departures from a nonparametric model. We note that if all true underlying models for the nuisance parameters include interactions, then in theory one can only construct estimators that are unbiased under the first and fourth specifications. To check sensitivity, we calculated the change in each estimator from estimates based on the first specification (saturated models), then scaling by the latter estimates.
@@
The results of the analysis can be found in Table 1. As might be expected, when all working models are saturated (except the exposure effect model), the estimates are in reasonable agreement. Fitting a linear model adjusted for radiation exposure, city of residence, sex and an interaction between the covariates yielded an exposure effext estimate of -0.32 (95% CI: -0.82, 0.21). A two-stage least squares assumption under the assumption that city was a valid instrument yielded an estimate of 3.48 (95% CI: 2.00, 4.96). Hence the estimates from our analysis tended to be arguably more plausible than from the alternative analyses, albeit with much wider confidence intervals. As one varies the model specifications, one can see the the multiply robust estimators produce stable estimates; in contrast, $\hat{\psi}_{g-Z}$, $\hat{\psi}_{g-S}$, $\hat{\psi}_{IPW}$ all appear to be highly sensitive to departures from fitting saturated models.
@@ 5 Discussion
There are also interesting connections with the proximal inference framework (Miao et al.,, 2018; Tchetgen Tchetgen et al.,, 2024). Here, a negative control outcome and exposure can together be leveraged to obtain nonparametric identification without a parallel trends assumption. In our case $\mathbf{Z}$ is a potentially invalid negative control exposure. Moreover, when the reference population is a negative control population, the negative control outcome is simply given by $(1-S)Y$, but because it is not measured in the target population, proximal inference cannot be directly applied. Hence our proposal can be viewed as an alternative to proximal inference where the negative control outcome is missing in the target population and the negative control exposure is invalid.
@@ A.1 Linear structural equation models
for $s=0,1$. If we first consider a conditional parallel trends or ‘additive equi-confounding’ condition (3), note that (3) can be shown to hold if $\zeta_{3,0}=\zeta_{3,1}$ and
$E(U|A=1,Z,S=1,\mathbf{C})-E(U|A=0,Z,S=1,\mathbf{C})$ (11)
@@
for $s=0,1$, a sufficient condition for (13) to hold is that $\zeta_{i,0}=\zeta_{i,1}$ for $i=0,...,3$ and $\upsilon_{j,0}=\upsilon_{j,1}$ for $j=0,1,2$. Hence in the presence of an unmeasured variable $U$, we require all associations encoded by the coefficients in (A.1) to transport across populations. We note that beyond model (A.1), Assumption 4 is weaker than full population exchangeability since it only restricts the additive interaction between $Z$ and $S$ with respect to the conditional mean of $Y^{0}$.
@@
  • Suppose that $\forall z$, $Y^{0,z}=Y^{0}$ (exclusion restriction). Then $\zeta_{1,0}=\zeta_{1,1}=0$ under the exclusion restriction. Hence Assumption 4 restricts the (conditional mean) association between $Y_{0}$ and $U$, as was done to justify (3), as well as the (conditional mean) association between $Z$ and $U$.
  • In this context, both additive equi-confounding (3) and Assumption 4 require that $\zeta_{3,0}=\zeta_{3,1}$. However, (3) is justified under (11) whereas Assumption 4 requires that $\upsilon_{1,0}=\upsilon_{1,1}$ as well as $\zeta_{1,0}=\zeta_{1,1}$.

### v6

@@ Abstract
Many proposals for the identification of causal effects require an instrumental variable that satisfies strong, untestable unconfoundedness and exclusion restriction assumptions. In this paper, we show how one can potentially identify causal effects under violations of these assumptions by harnessing a negative control population or outcome. This strategy allows one to leverage sub-populations for whom the exposure is degenerate, and requires that the instrument-outcome association satisfies a certain parallel trend condition. We develop the semiparametric efficiency theory for a general instrumental variable model, and obtain a multiply robust, locally efficient estimator of the average treatment effect in the treated. The utility of the estimators is demonstrated in simulation studies and an analysis of the Life Span Study.
@@ 1 Introduction
Alternatively, a negative control outcome is a variable that cannot plausibly be caused by the exposure. Under the core instrumental variable assumptions, the instrument would then not be associated with the negative control outcome. For example, one might choose a clinical outcome unlikely to be affected by the medication. There is an existing literature on using negative control outcomes as a means to detect and remove bias in an analysis based on confounding adjustment (Lipsitch et al.,, 2010; Sofer et al.,, 2016); however, they have been used less for assessing the instrumental variable assumptions. Davies et al., (2017) describe how negative control outcomes can be used to compare the plausibility of instrumental variable approaches with analyses that assume no unmeasured confounding; Sanderson et al., (2021) use these variables to check for violations of unconfoundedness due to population stratification in Mendelian Randomisation studies.
We go beyond falsification testing and propose conditions for identification and estimation of effects with an invalid instrument variable. Our results rely on a condition on the stability of the association between the instrument and outcome, which is closely related to the parallel trends assumption in the difference-in-difference literature. Intuitively, our assumption states that the instrument-outcome association that arises due to violations of IV.2 and IV.3 in the naïve instrumental variable analysis must equal the instrument-outcome association in a reference population. If our assumption holds, the potential instrument does not necessarily have to satisfy the conditions IV.2 and IV.3, but still must be predictive of the exposure, therefore satisfying IV.1. A special feature of our framework, relative to other negative control strategies, is that it allows one to leverage sub-populations who were not eligible to be exposed. We hope that our methodology may be generally useful in epidemiologic applications where such sub-populations arise.
@@ 2.1 Bias detection using reference populations
Reference population: Let $\mathcal{C}_{0}$ ($\mathcal{C}_{1}$) denote the support of $\mathbf{C}$ in individuals with $S=0$ ($S=1$). Then $\forall z\in\{0,1\}$ and $\forall\mathbf{c}\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$,
@@
A negative control population where the exposure is not necessarily degenerate but is hypothesised to have no effect, could also form a reference population. However, the concept of a reference population allows for the possibility that the treatment causally impacts the outcome in individuals with $S=0$, if these individuals were not excluded from access to treatment. In general, it is advantageous to choose a reference population that is similar in important baseline characteristics to the $S=1$ population. This is because to eventually identify a causal effect, we will need to transport associations across from the reference to the $S=1$ population. We emphasise that it is not obvious how other negative control-type designs can generally leverage sub-populations for whom treatment is degenerate. These designs usually involve contrasts between levels of $A$ that are expected to be zero when there is no bias. Hence some variation in exposure is typically required (Lipsitch et al.,, 2010).
$Z$$A$$C$$Y$$U$(a) Population of interest ($S=1$)
@@
See B.1 of the appendix for a proof. This statement provides the basis of a valid test of IV.2 and IV.3 in the reference population. To draw conclusions about the population of interest, we would furthermore need to assume that $\forall z\in\{0,1\}$ and $\forall\mathbf{c}\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$,
@@
In other words, $Z$ is an instrument in $S=0$ if it is an instrument in $S=1$. Then (1) can be used as the basis of a valid falsification test for the IV assumptions in the population of interest. If we strengthen (2.1) to an ‘iff’ statement and Proposition 2.0.1 to an ‘iff’ result (under a faithfulness-type condition), then we can construct a consistent test. We note that the proposition allows for the distribution of $\mathbf{C}$ to differ between populations, but if the support of $\mathbf{C}$ differs between the standard and reference populations, one may not be able to draw conclusions for all individuals with $S=1$.
@@ 2.2 Point identification
We will focus on identification of the conditional average treatment effect on the treated in the population of interest:
@@
We note first that Assumption 2 is standard in causal inference. Assumption 3 demands that $Z$ is predictive of the exposure; although this assumption is similar to IV.1, $Z$ is not required to satisfy either IV.2 or IV.3. In particular, our assumptions also allow for $Z$ to be a standard confounder with a direct effect on the outcome (violating the exclusion restriction). When IV.2 and IV.3 hold in both populations, we note that Assumption 4 then automatically also holds. This is because one can show that under the IV conditions, $E(Y^{0}|Z=1,S=s,\mathbf{C}=\mathbf{c})-E(Y^{0}|Z=0,S=s,\mathbf{C}=\mathbf{c})=0$ $\forall\mathbf{c}\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$ and $s=0,1$. Assumption 4 requires that the (conditional) additive association between the potential outcome $Y_{0}$ and $Z$ transports between the negative control population and the and population of interest. It is not generally testable. We discuss the plausibility of this assumption below, which is weaker than the full population exchangeability assumption made in the transportability literature (Dahabreh et al.,, 2019).
@@
Our assumptions are stated conditional on covariates $\mathbf{C}$. To ensure Assumption 4 holds, we require that $\mathbf{C}$ includes all effect modifiers of the additive association between $Z$ and $Y^{0}$ that differ in distribution between populations. In terms of design implications, this suggests choosing a reference population that is similar to the population of interest in terms of characteristics that may modify the association between $Z$ and $Y^{0}$. As argued above, it is also advantageous to incorporate measured confounders for the effect of $Z$ on $Y$, as well as for the effect of $A$ on $Y$, since this may weaken the restrictions on $U$.
@@ 2.3 Negative control outcomes
One can also harness measurements of a negative control outcome in order to generate a reference population. Suppose that there is now only data on individuals from the population of interest, such that $S=1$ for everyone (we omit this from the conditioning statement to simplify notation). Further, suppose that one collects a random variable $W$ in the population of interest that satisfies the following assumption
@@ 3.1 Semiparametric Theory
$\Omega\equiv\{\phi(\mathbf{Z},S,\mathbf{C}):E\{\phi(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}=E\{\phi(\mathbf{Z},S,\mathbf{C})|S,\mathbf{C}\}=0\}$
@@
$-E^{\ddagger}\{r_{0}(\mathbf{Z},S,\mathbf{C})|S,\mathbf{C}\}+E^{\ddagger}\{r_{0}(\mathbf{Z},S,\mathbf{C})|\mathbf{C}\}]:\textrm{$r_{0}(\mathbf{Z},S,\mathbf{C})$ unrestricted}\bigg{\}}$
@@ 3.2 De-biased machine learning and multiply robust estimation
The estimation strategies previously described are not generally feasible, because the estimating equations for $\psi^{\dagger}$ involve nuisance parameters that are typically unknown. One approach to deal with this is to plug in estimates obtained from nonparametric estimators or flexible statistical learning methods (Chernozhukov et al.,, 2018). In A.6 of the appendix, we therefore use Theorem 3.1 to construct cross-fit de-biased machine learning-based estimators of $\psi^{\dagger}$, and describe sandwich estimators of the asymptotic variance. Nevertheless, to develop a more nuanced understanding of the different bias properties of RAL estimators of $\psi^{\dagger}$, we will instead consider a scenario where parametric working models are used for the conditional expectations/densities that arise in estimating the causal effect of interest. We hence conceptualise bias as potentially arising from parametric model misspecification. The parametric approach will also suggest how nuisance functions should be estimated nonparametrically for optimal performance in terms of their convergence rates.
@@
  • $\mathcal{M}_{3}$: $f(S|\mathbf{Z}=\mathbf{0},\mathbf{C})=f(S|\mathbf{Z}=\mathbf{0},\mathbf{C};\alpha^{\dagger})$, $OR(\mathbf{Z},S,\mathbf{C})=OR(\mathbf{Z},S,\mathbf{C};\rho^{\dagger})$, and $b_{1}(\mathbf{C})=b_{1}(\mathbf{C};\theta^{\dagger}_{1})$.
  • $\mathcal{M}_{4}$: $f(\mathbf{Z}|S=0,\mathbf{C})=f(\mathbf{Z}|S=0,\mathbf{C};\tau^{\dagger})$, $f(S|\mathbf{Z}=\mathbf{0},\mathbf{C})=f(S|\mathbf{Z}=\mathbf{0},\mathbf{C};\alpha^{\dagger})$ and $OR(\mathbf{Z},S,\mathbf{C})=OR(\mathbf{Z},S,\mathbf{C};\rho^{\dagger})$.
At one extreme, $\mathcal{M}_{1}$ requires a correct conditional mean model for outcome after removing the treatment effect: $E\{Y-\beta(\mathbf{A,C})S|\mathbf{Z},S,\mathbf{C}\}$. At the other, $\mathcal{M}_{4}$ requires a correct model for the joint conditional density $f(\mathbf{Z},S|\mathbf{C})$. However, we allow for certain combinations of restrictions on these laws. If our model for $f(S|\mathbf{Z},\mathbf{C})$ is correct, then under $\mathcal{M}_{3}$, assumptions on the impact of $\mathbf{Z}$ on the conditional mean of $Y$ given $\mathbf{C}$ in the reference population can be relaxed.
@@ 4 Data analysis
Varying the fitted models in this way enabled us to check the robustness of the estimators $\hat{\psi}_{TSLS}$, $\hat{\psi}_{g-Z}$, $\hat{\psi}_{g-S}$, $\hat{\psi}_{IPW}$, $\hat{\psi}_{MR}$ and $\hat{\psi}_{MR-eff}$ to departures from a less restrictive model. We note that if all true underlying models for the nuisance parameters include interactions, then in theory one can only construct estimators that are unbiased under the first and fourth specifications. To check sensitivity, we calculated the change in each estimator from estimates based on the first specification (saturated models), then scaling by the latter estimates.
@@
The results of the analysis can be found in Table 1. As might be expected, when all working models are saturated (except the exposure effect model), the estimates are in reasonable agreement. Fitting a linear model adjusted for radiation exposure, city of residence, sex and an interaction between the covariates yielded an exposure effext estimate of -0.32 (95% CI: -0.82, 0.21). A two-stage least squares assumption under the assumption that city was a valid instrument yielded an estimate of 3.48 (95% CI: 2.00, 4.96). Hence the estimates from our analysis tended to be arguably more plausible than from the alternative analyses, albeit with much wider confidence intervals. As one varies the model specifications, one can see the multiply robust estimators produce stable estimates; in contrast, $\hat{\psi}_{g-Z}$, $\hat{\psi}_{g-S}$, $\hat{\psi}_{IPW}$ all appear to be highly sensitive to departures from fitting saturated models.
@@ 5 Discussion
There are also interesting connections with the proximal inference framework (Miao et al.,, 2018; Tchetgen Tchetgen et al.,, 2024). Here, a negative control outcome and exposure can together be leveraged to obtain nonparametric identification without a parallel trends assumption. In our case $\mathbf{Z}$ is a potentially invalid negative control exposure. Moreover, when the reference population is a negative control population, the negative control outcome is simply given by $(1-S)Y$, but because it is not measured in the population of interest, proximal inference cannot be directly applied. Hence our proposal can be viewed as an alternative to proximal inference where the negative control outcome is missing in the population of interest and the negative control exposure is invalid.
@@ A.1 Linear structural equation models
for $s=0,1$. For simplicity, we let $U$ be a single random variable. If we first consider a conditional parallel trends or ‘additive equi-confounding’ condition (3), note that (3) can be shown to hold if $\zeta_{3,0}=\zeta_{3,1}$ and
$E(U|A=1,Z,S=1,\mathbf{C})-E(U|A=0,Z,S=1,\mathbf{C})$
@@
for $s=0,1$, a sufficient condition for (11) to hold is that $\zeta_{i,0}=\zeta_{i,1}$ for $i=0,1,2,3$ and $\upsilon_{j,0}=\upsilon_{j,1}$ for $j=0,1,2$. Hence in the presence of an unmeasured variable $U$, we require all associations encoded by the coefficients in (A.1) to transport across populations. We note that beyond model (A.1), Assumption 4 is weaker than full population exchangeability since it only restricts the additive interaction between $Z$ and $S$ with respect to the conditional mean of $Y^{0}$.
@@
  • Suppose that $\forall z$, $Y^{0,z}=Y^{0}$ (exclusion restriction). Then $\zeta_{1,0}=\zeta_{1,1}=0$ under the exclusion restriction. Hence Assumption 4 restricts the (conditional mean) association between $Y^{0}$ and $U$, as was done to justify (3), as well as the (conditional mean) association between $Z$ and $U$.
  • In this context, both additive equi-confounding (3) and Assumption 4 require that $\zeta_{3,0}=\zeta_{3,1}$. However, (3) is justified under (A.1) whereas Assumption 4 requires that $\upsilon_{1,0}=\upsilon_{1,1}$ as well as $\zeta_{1,0}=\zeta_{1,1}$.